OpenStudy (anonymous):

Write an equation of the line that passes through the point (2, 3) and is perpendicular to the line x = -1

OpenStudy (anonymous):

x=-1 0=x+1 0y=x+1 its slope =1/0=infinity slope of line perpendicular to it=-1/infinity=0 required eq. of line is y-y1=m(x-x1)

OpenStudy (anonymous):